Hello! I need to solve this equation by completing the square:
5x^2+15x-1 =0
I know that the answer is -15-7√5/10, -15+7√5/10, I just don't know how to show all of the work. I know that I divide everything by 5, and you have a 3x. Then, take half of then and then square it (leaving it as a fraction).
This is due in the morning and I'm stuck. I'd appreciate help! Thanks!

Answer by Alan3354(69443)

5x^2+15x-1 =0
5x^2+15x = 1
x^2 + 3x = 1/5
x^2 + 3x + (3/2)^2 = 1/5 + (3/2)^2
x^2 + 3x + 9/4 = 1/5 + 9/4 = 49/20
etc